Airplane travels 700miles agains wind in 4 hours. And makes a return trip with the same wind in 1 hour. Find speed of wind.
Answer by checkley77(12844)

700/4=175 mph. is the speed against the wind.
700/1=700 mph. is the speed with the wind.
(700-175)/2=525/2=262.5 mph is the speed of the wind.
700-262.5=437.5 mph. is the speed of the plane in still air.
175+262.5=437.5 ditto.